OpenStudy (anonymous):

Charlotte has two job offers as a car sales person. The first job will pay her $600 a month plus 2% on the price of her car sales. The second job will pay $1000 a month plus 1% on the price of her sales. How much must she sell each month to make the same income from the two jobs? $20,000 $30,000 $40,000 $50,000

OpenStudy (anonymous):

Well, the amount she would make from the first job can be expressed as\[m=600+0.02 \times s\]where m is the money she makes and s is the price of her sales. The second job would be\[m=1000+0.01 \times s\]Just set them equal and then solve for s:\[600+0.02 \times s=1000+0.01 \times s\]